OpenStudy (anonymous):

Find two consecutive even integers such that twice the smaller is 16 more than the larger.

OpenStudy (anonymous):

let them be 2x & 2x +2 2(2x) = 2x +2 + 16 4x = 2x + 18 4x - 2x = 18 2x = 18 x = 9 the numbers are 2x & 2x + 2 substitute x and find it

OpenStudy (anonymous):

18 and 20. Okay cool. I think I get it now.